OpenStudy (anonymous):

if 3a+1=x and 3a=y,express y in terms of x

OpenStudy (texaschic101):

3a = y -- sub 3a in for y in the other equation 3a + 1 = x y + 1 = x y = x - 1 or x = y + 1

OpenStudy (texaschic101):

in terms of x, I guess it would be : x = y + 1

OpenStudy (anonymous):

in terms of x it would be y=x-1. The question asked you to express y. That normally means "y= (some expression)"